# Active learning of neural population dynamics using two-photon holographic optogenetics

### v2

@@ 4.1 Constrained Nuclear Norm Estimator under Non-Isotropic Measurements
for any $M\in\mathbb{R}^{d_{1}\times d_{2}}$. We call $P_{\parallel}$ the projection onto the tangent space of $\Theta_{\star}$. Note that the dimension of the range of $P_{\parallel}$ is equal to just $r(d_{1}+d_{2})-r^{2}\ll d_{1}d_{2}$. We are now ready to state our main result on the estimation error of $\widehat{\Theta}$, for $\widehat{\Theta}$ as defined in (4.2).
@@
Define $\mu:=\|(\Phi^{*}\Phi)^{1/2}((P_{\parallel}\Phi^{*}\Phi P_{\parallel})^{\dagger})^{1/2}\|_{\mathrm{op}}$. Then with probability at least $1-2\delta$:
@@ Appendix A Proof of Theorem 1
$\langle\Phi(\Delta),v\rangle=\langle\Delta,\sum_{n=1}^{N}X_{n}v_{n}\rangle=:\langle\Delta,\Phi^{*}(v)\rangle$
where $\Phi^{*}$ denotes the adjoint of $\Phi$. We also recognize that the operator $\Phi^{*}\Phi:\mathbb{R}^{d_{1}\times d_{2}}\rightarrow\mathbb{R}^{d_{1}\times d_{2}}$ is also linear, defined as $\Phi^{*}\Phi(M)=\Phi^{*}(\{\langle X_{n},M\rangle\}_{n})=\sum_{n=1}^{N}X_{n}\langle X_{n},M\rangle$. Consequently $(\Phi^{*}\Phi)^{1/2}$ is well-defined and is the same operator as $\Phi^{*}\Phi$ after taking the square root of its eigenvalues.
@@
$=\max_{\Delta:\|P_{\perp}(\Delta)\|_{*}\leq t}2\langle\sum_{n=1}^{N}\eta_{n}X_{n},P_{\perp}(\Delta)\rangle$
$=\max_{\Delta:\|P_{\perp}(\Delta)\|_{*}\leq t}2\langle P_{\perp}(\sum_{n=1}^{N}\eta_{n}X_{n}),P_{\perp}(\Delta)\rangle$
$\leq 2t\|P_{\perp}(\sum_{n=1}^{N}\eta_{n}X_{n})\|_{\mathrm{op}}$
@@
$\sum_{n=1}^{N}\eta_{n}X_{n}=\text{mat}(\sum_{n=1}^{N}\eta_{n}\text{vec}(X_{i}))=\text{mat}((\sum_{n=1}^{N}\text{vec}(X_{n})\text{vec}(X_{n})^{\top})^{1/2}\text{vec}(\eta^{\prime}))=:(\Phi^{*}\Phi)^{1/2}(\eta)$
@@
$\|P_{\perp}(\sum_{n=1}^{N}\eta_{n}X_{n})\|_{\mathrm{op}}$ $=\|P_{\perp}(\text{mat}((\sum_{n=1}^{N}\text{vec}(X_{n})\text{vec}(X_{n})^{\top})^{1/2}\text{vec}(\eta^{\prime})))\|_{\mathrm{op}}$
@@
Define $\mu:=\|(\Phi^{*}\Phi)^{1/2}((P_{\parallel}\Phi^{*}\Phi P_{\parallel})^{\dagger})^{1/2}\|_{\mathrm{op}}$. Then
@@
To aid in readability, we make a number of notational modifications. First, we drop parentheses so that $\Phi(P_{\parallel}(\Delta))$ is just notated as $\Phi P_{\parallel}\Delta$. Second, we define $M^{\dagger/2}:=(M^{\dagger})^{1/2}$ where $M^{\dagger}$ is the pseudoinverse. If $\Phi^{*}\Phi$ is invertible restricted to the range of $P_{\parallel}$, then $P_{\parallel}\Delta=(P_{\parallel}\Phi^{*}\Phi P_{\parallel})^{\dagger/2}(\Phi^{*}\Phi)^{1/2}P_{\parallel}\Delta$ for all $\Delta$. Thus,
@@
$=|\langle P_{\parallel}\Delta,(\Phi^{*}\Phi)P_{\perp}\Delta\rangle|$
$=|\langle(\Phi^{*}\Phi)^{1/2}P_{\parallel}\Delta,(\Phi^{*}\Phi)^{1/2}P_{\perp}\Delta\rangle|$
$=|\langle(\Phi^{*}\Phi)^{1/2}P_{\parallel}(P_{\parallel}\Phi^{*}\Phi P_{\parallel})^{\dagger/2}(\Phi^{*}\Phi)^{1/2}P_{\parallel}\Delta,(\Phi^{*}\Phi)^{1/2}P_{\perp}\Delta\rangle|$
$\leq\|(\Phi^{*}\Phi)^{1/2}(P_{\parallel}\Phi^{*}\Phi P_{\parallel})^{\dagger/2}(\Phi^{*}\Phi)^{1/2}P_{\parallel}\Delta\|_{\mathrm{F}}\,\|(\Phi^{*}\Phi)^{1/2}P_{\perp}\Delta\|_{\mathrm{F}}$
$\leq\|(\Phi^{*}\Phi)^{1/2}(P_{\parallel}\Phi^{*}\Phi P_{\parallel})^{\dagger/2}\|_{\mathrm{op}}\,\|(\Phi^{*}\Phi)^{1/2}P_{\parallel}\Delta\|_{\mathrm{F}}\,\|(\Phi^{*}\Phi)^{1/2}P_{\perp}\Delta\|_{\mathrm{F}}$
@@
Thus, if $\mu:=\|(\Phi^{*}\Phi)^{1/2}(P_{\parallel}\Phi^{*}\Phi P_{\parallel})^{\dagger/2}\|_{\mathrm{op}}$ then
@@
$\geq\|\Phi(P_{\parallel}(\Delta))\|_{2}^{2}+\|\Phi(P_{\perp}(\Delta))\|_{2}^{2}-2\mu\|(\Phi^{*}\Phi)^{1/2}P_{\parallel}\Delta\|_{\mathrm{F}}\,\|(\Phi^{*}\Phi)^{1/2}P_{\perp}\Delta\|_{\mathrm{F}}$

### v4

@@ 4.1 Constrained Nuclear Norm Estimator under Non-Isotropic Measurements
for any $M\in\mathbb{R}^{d_{1}\times d_{2}}$. We call $P_{\parallel}$ the projection onto the tangent space of $\Theta_{\star}$. Note that the dimension of the range of $P_{\parallel}$ is equal to just $r(d_{1}+d_{2})-r^{2}\ll d_{1}d_{2}$. When convenient, we drop parentheses so that $\Phi(P_{\parallel}(\Delta))$ is just notated as $\Phi P_{\parallel}\Delta$. Since $\Phi(P_{\parallel}(\cdot))$ is a linear operator from $\mathbb{R}^{d_{1}\times d_{2}}\rightarrow\mathbb{R}^{N}$, the adjoint $(\Phi P_{\parallel})^{*}=P_{\parallel}^{*}\Phi^{*}=P_{\parallel}\Phi^{*}$ and Moore-Penrose inverse $(\Phi P_{\parallel})^{\dagger}$ are well-defined. We are now ready to state our main result on the estimation error of $\widehat{\Theta}$, for $\widehat{\Theta}$ as defined in (4.2).
@@
Define $\mu:=\|(P_{\parallel}\Phi^{*})^{\dagger}P_{\parallel}\Phi^{*}\Phi P_{\perp}(\Phi P_{\perp})^{\dagger}\|_{\mathrm{op}}$. Then with probability at least $1-2\delta$:
@@ Appendix A Proof of Theorem 1
$\langle\Phi(\Delta),v\rangle=\sum_{n=1}^{N}\langle\Delta,\varphi_{n}\rangle v_{n}=\langle\Delta,\sum_{n=1}^{N}\varphi_{n}v_{n}\rangle=:\langle\Delta,\Phi^{*}(v)\rangle$
where $\Phi^{*}$ denotes the adjoint of $\Phi$. We also recognize that the operator $\Phi^{*}\Phi:\mathbb{R}^{d_{1}\times d_{2}}\rightarrow\mathbb{R}^{d_{1}\times d_{2}}$ is also linear, defined as $\Phi^{*}\Phi(M)=\Phi^{*}(\{\langle\varphi_{n},M\rangle\}_{n})=\sum_{n=1}^{N}\varphi_{n}\langle\varphi_{n},M\rangle$. Consequently $(\Phi^{*}\Phi)^{1/2}$ is well-defined and is the same operator as $\Phi^{*}\Phi$ after taking the square root of its eigenvalues.
@@
$=\max_{\Delta:\|P_{\perp}(\Delta)\|_{*}\leq t}2\langle\sum_{n=1}^{N}\eta_{n}\varphi_{n},P_{\perp}(\Delta)\rangle$
$=\max_{\Delta:\|P_{\perp}(\Delta)\|_{*}\leq t}2\langle P_{\perp}(\sum_{n=1}^{N}\eta_{n}\varphi_{n}),P_{\perp}(\Delta)\rangle$
$\leq 2t\|P_{\perp}(\sum_{n=1}^{N}\eta_{n}\varphi_{n})\|_{\mathrm{op}}$
@@
$\sum_{n=1}^{N}\eta_{n}\varphi_{n}=\text{mat}(\sum_{n=1}^{N}\eta_{n}\text{vec}(\varphi_{n}))=\text{mat}((\sum_{n=1}^{N}\text{vec}(\varphi_{n})\text{vec}(\varphi_{n})^{\top})^{1/2}\text{vec}(\eta^{\prime}))=:(\Phi^{*}\Phi)^{1/2}(\eta)$
@@
$\|P_{\perp}(\sum_{n=1}^{N}\eta_{n}\varphi_{n})\|_{\mathrm{op}}$ $=\|P_{\perp}(\text{mat}((\sum_{n=1}^{N}\text{vec}(\varphi_{n})\text{vec}(\varphi_{n})^{\top})^{1/2}\text{vec}(\eta^{\prime})))\|_{\mathrm{op}}$
@@
Define $\mu:=\|(P_{\parallel}\Phi^{*})^{\dagger}P_{\parallel}\Phi^{*}\Phi P_{\perp}(\Phi P_{\perp})^{\dagger}\|_{\text{op}}$. Then
@@
To aid in readability, we make a number of notational modifications. First, we drop parentheses so that $\Phi(P_{\parallel}(\Delta))$ is just notated as $\Phi P_{\parallel}\Delta$. Noting that $\Phi(P_{\parallel}(\cdot))$ is a linear operator from $\mathbb{R}^{d_{1}\times d_{2}}\rightarrow\mathbb{R}^{N}$ it can be thought of as a matrix operating on the vectorized $\Delta$. Thus, the adjoint $(\Phi P_{\parallel})^{*}=P_{\parallel}^{*}\Phi^{*}=P_{\parallel}\Phi^{*}$ and Moore-Penrose inverse $(\Phi P_{\parallel})^{\dagger}$ are well-defined. Then
@@
$=|\langle\Phi P_{\parallel}(\Phi P_{\parallel})^{\dagger}\Phi P_{\parallel}\Delta,\Phi P_{\perp}(\Phi P_{\perp})^{\dagger}\Phi P_{\perp}\Delta\rangle|$
$=|\langle\Phi P_{\parallel}\Delta,(P_{\parallel}\Phi^{*})^{\dagger}P_{\parallel}\Phi^{*}\Phi P_{\perp}(\Phi P_{\perp})^{\dagger}\Phi P_{\perp}\Delta\rangle|$
$\leq\|\Phi P_{\parallel}\Delta\|_{2}\,\|(P_{\parallel}\Phi^{*})^{\dagger}P_{\parallel}\Phi^{*}\Phi P_{\perp}(\Phi P_{\perp})^{\dagger}\Phi P_{\perp}\Delta\|_{2}$
$\leq\|\Phi P_{\parallel}\Delta\|_{2}\,\|\Phi P_{\perp}\Delta\|_{2}\,\|(P_{\parallel}\Phi^{*})^{\dagger}P_{\parallel}\Phi^{*}\Phi P_{\perp}(\Phi P_{\perp})^{\dagger}\|_{\text{op}}$
@@
Thus, if $\mu:=\|(P_{\parallel}\Phi^{*})^{\dagger}P_{\parallel}\Phi^{*}\Phi P_{\perp}(\Phi P_{\perp})^{\dagger}\|_{\text{op}}$ then
@@
$\geq\|\Phi(P_{\parallel}(\Delta))\|_{2}^{2}+\|\Phi(P_{\perp}(\Delta))\|_{2}^{2}-2\mu\|\Phi(P_{\parallel}(\Delta))\|_{2}\|\Phi(P_{\perp}(\Delta))\|_{2}$